OpenStudy (korosh23):

**Physics 12 Question** Topic: Electrostatics Calculate the electric field strength mid-way between 4.50 x 10^-6 C charged object and a -4.50 x 10^-6 C charged object if the two charged objects are 5.00 x 10^-1 m apart. Answer: 1.296 x 10^6 N/C Here is the answer. Show me how to get that. Thank you.

OpenStudy (faiqraees):

Are you sure the answer and the units used are correct?

OpenStudy (aravindg):

Find the electric field produced at the midpoint due to each charge separately. Use the formula E=kq/r^2 .Now add these 2 electric fields so that you get net electric field. Since charges are identical and equidistant, it will be 2E. Plug in and get your answer. I got the same answer as yours when I did it.

OpenStudy (michele_laino):

hint: the situation can be represented as below:  the magnitude of the resultant electric field at midpoint M, is: \[E = k\frac{Q}{{{{\left( {d/2} \right)}^2}}} + k\frac{Q}{{{{\left( {d/2} \right)}^2}}} = 2k\frac{Q}{{{{\left( {d/2} \right)}^2}}}\]